OpenStudy (anonymous):

Solve the following system of equations: 2x - y + z = -3 2x + 2y + 3z = 2 3x - 3y - z = -4

OpenStudy (anonymous):

Do you know how to do Matrices?

OpenStudy (anonymous):

no @ChrisTopher987654321

OpenStudy (anonymous):

I'd go on to the Khan academy website and learn that. Because without it you wouldn't have a structured way of doing it.

OpenStudy (shinalcantara):

you may apply elimination and substitution method

OpenStudy (anonymous):

I tried that andkept on messingup @shinalcantara

OpenStudy (anonymous):

Here the knanacademy video that will help you https://www.khanacademy.org/math/algebra/systems-of-eq-and-ineq/fancier-systems/v/systems-of-three-variables

OpenStudy (anonymous):

Here's*

OpenStudy (anonymous):

tnx @faith2012

OpenStudy (anonymous):

You're welcome! I hope it'll help you.

OpenStudy (shinalcantara):

as long as the number of unknowns is equal to the number of equations, you'll be able to get the values using elimination and substitution method.. though you may be able to answer this just by inputting all the coefficients to the calculator using MODE 5, 2 of your fx-991ES PLUS calculator, you have to know the basic.. :)

OpenStudy (shinalcantara):

2x-y+z=-3 --- eq'n 1 2x+2y+3z=2 ----eq'n 2 3x-3y-z=-4 ------eq'n 3 add eq'ns 1& 3 you'll get 5x-4y=-7 -- ---let this be eq'n 4 rearrange eq'n 3 such that 'z' would have a value in terms of x & y z = 4+3x-3y ------eq'n 5 substitute eq'n 5 to eq'n 2 2x+2y+3(4+3x-3y)=2 2x+2y+12+9x-9y=2 11x-7y=-10 -----eq'n 6 equate eq'ns 4 & 6 and multiply eqn 4 with 7 and eqn 6 with 4 you'll have: (5x-4y=-7)*7 ----- 35x-28y=-49 ---eq'n 7 (11x-7y=-10)*4 44x-28y=-40 ---eq'n 8 subtract eqn 7 from eqn 8 9x=9 x=1 substitute to eq'n 4 5x-4y=-7 5(1)-4y=-7 4y=12 y=3 substitute to eqn 5 z=4+3(1)-3(3) z=-2